# Synthetic Data Generation in Cybersecurity: A Comparative Analysis

### v1

@@ Abstract
Synthetic data generation faces significant challenges in accurately replicating real data, particularly with tabular data, where achieving high fidelity and utility is critical. While numerous methods have been developed, the most effective approach for creating high-quality synthetic data for network traffic security remains to be seen. This study conducts a comprehensive comparative analysis of non-AI, conventional AI, and generative AI techniques for synthetic tabular data generation using two widely recognized cybersecurity datasets: NSL-KDD and CICIDS-2017. Particular emphasis was placed on prominent GAN models for tabular data generation, including CTGAN, CopulaGAN, GANBLR++, and CastGAN. The results indicate that GAN-based methods, particularly CTGAN and CopulaGAN, outperform non-AI and conventional AI approaches in terms of fidelity and utility. To the best of our knowledge, this research contributes to the field by offering the first comparative evaluation of these methods specifically for cybersecurity network traffic data, filling a critical gap in the literature. It also introduces mutual information for feature selection, further enhancing the quality of the generated synthetic data. These findings provide valuable guidance for researchers seeking the most suitable synthetic data generation method in cybersecurity applications.
@@
Generative Adversarial Networks (GANs), Synthetic tabular Data, Cybersecurity, NSL-KDD, CICIDS, Mutual information, Feature selection
@@ I Introduction
In the rapidly evolving field of cybersecurity, IDS plays a crucial role in identifying and mitigating threats within network environments. These systems rely heavily on high-quality data to detect patterns indicative of malicious activity [1]. However, acquiring real-world cybersecurity data poses significant challenges, primarily due to privacy concerns and time constraints [2]. Researchers have turned to synthetic data generation as a viable alternative to address this. Synthetic data, when adequately generated, can mimic the characteristics of real-world data while avoiding the associated ethical and legal issues, making it a valuable resource for training and evaluating IDS [3].
Despite the growing interest in synthetic data generation, producing synthetic datasets that maintain the utility, fidelity, and data integrity required for practical cybersecurity applications remains a significant challenge [4]. Utility refers to the ability of synthetic data to support the same analytical tasks as real data [5]. At the same time, fidelity ensures that the synthetic data accurately reflects the statistical properties of the original data[6]. Data integrity, on the other hand, focuses on the accuracy and reliability of data, ensuring that it remains unaltered and consistent [7, 8]. Follow these criteria to avoid synthetic data that is either unrealistic or similar to real data, rendering it less valuable or risky.
GANs have shown great potential for generating synthetic tabular data for cybersecurity purposes. Various GAN architectures like VanillaGAN[9], WGAN[10], and WGAN-GP[11], as well as models specifically designed for tabular data generation such as CTGAN[12], CopulaGAN [13], and GANBLR [14], have been successfully used to create synthetic IDS data [15]. However, the effectiveness of GANs for generating tabular data specific to cybersecurity, such as network flow data used in IDS, has yet to be thoroughly benchmarked against other tabular synthetic data generation methods: Artificial intelligence (AI) based and non-AI based.
I-A Synthetic data for cybersecurity
@@
    Data privacy: Sharing real-world cybersecurity data involves a risk of sharing sensitive information, which hinders collaboration and research [16]. Thus, the synthetic data resolves this issue, and valuable data can be shared without compromising privacy [17].
@@
    Data scarcity: The real-world cyber security data is often imbalanced, with normal traffic making up a higher percentage than attacks or malicious traffic. This causes a problem of data scarcity. Synthetic data in this scenario can help balance both instances and improve cyber security tasks, such as efficient, unbiased intrusion detection [1].
@@
    Diversity of scenarios: Cybersecurity threats are becoming more dynamic and diverse, but real cybersecurity datasets often need more diversity [18]. In this scenario, synthetic data helps simulate various uncommon attacks and allows for more comprehensive testing of cybersecurity tools [19].
@@
    Enhancing AI models: When training cyber security models such as Intrusion detection systems (IDS), a large, high-quality dataset is essential for effective detection. Therefore, synthetic data plays a crucial role in training these models on a diverse and extensive dataset to enhance the robustness of AI models [20, 21].
I-B Synthetic data challenges in cybersecurity
@@
    Data realism: One of the primary challenges in generating synthetic data is ensuring its realism compared to real-world data [22]. The effectiveness of AI models trained on synthetic data depends on how accurately the data represents real cyber threats and network conditions[23]. Many generative models are still in the process of achieving data realism by increasing their fidelity and utility to mimic actual trends [24].
@@
    Lack of variability: Some generative models are explicitly designed to generate a diverse set of network instances and cyber threats. However, if synthetic data is not generated with sufficient variability, it may miss capturing the full spectrum of possible attack vectors in cybersecurity [25, 26]. For instance, the Wasserstein GAN has solved the problem faced by the vanilla GAN of mode collapse, thus ensuring that the generated outcome is not limited to a few instances [10].
@@
    Class imbalance: Cybersecurity often suffers from class imbalance, where regular traffic is much larger than cyberattacks [27]. Synthetic data generation methods address this problem by balancing malicious and benign traffic and attacks within the cybersecurity dataset for more robust intrusion detection [28].
@@
    Privacy concerns: Synthetic data attempts to conceal confidential information from real-world sensitive data to address privacy and security concerns [29]. However, it is not entirely free from these challenges. Ensuring that sensitive information cannot be traced back from the synthetic data is crucial and remains a concern.
  5. 5.
    Standardization: More standardization is needed for generating, validating, and integrating synthetic data into cybersecurity practices. This makes it challenging to ensure the reliability and consistency of the synthetic datasets across different cybersecurity applications [30, 31].
  6. 6.
    Perfomace evaluation: For evaluating synthetic data, it is crucial to note that achieving a high similarity score does not guarantee its real-world applicability. Similarly, obtaining high accuracy in synthetic cybersecurity datasets for any ML/DL model does not ensure compatibility with real-world datasets [24, 32].
@@ I-C Motivation
This research addresses critical challenges in strengthening IDS defenses amidst a rapidly escalating cyber threat landscape. IDS are vital for securing networks by detecting and mitigating intrusions, but a lack of high-quality, real-world data for training and refinement hampers their effectiveness. Obtaining such data is challenging due to privacy concerns, ethical issues, and evolving network behaviors. Synthetic data generation using techniques like GANs offers a promising solution by creating realistic, high-fidelity data that mimics network activities without violating privacy or legal constraints. This research is crucial to improving IDS capabilities, enabling them to adapt to modern cyber adversaries’ sophisticated tactics. This study aims to enhance IDS robustness and reliability by benchmarking different synthetic data generation methods, contributing to more secure cybersecurity frameworks.
This research fills a gap in cybersecurity by offering a comprehensive comparison of synthetic data generation methods for enhancing IDS. While synthetic data’s potential has been recognized, there is limited clarity on which methods—traditional, AI-based, or GAN-based—provide the highest fidelity and utility for IDS. Existing literature has explored various techniques but lacks a focused comparison of their effectiveness in generating network traffic data for IDS. This study bridges the gap by systematically evaluating different synthetic data generation techniques, guiding the development of more effective IDS and contributing to more robust, more adaptable cybersecurity defenses.
@@ I-D Objectives
The primary objective of this study is to rigorously evaluate and compare the performance of various synthetic data generation methods for cybersecurity, focusing on their utility, fidelity, data integrity, and ability to handle class imbalance. To achieve this, the research addresses the following key questions:
@@
    Which methods produce synthetic data that best preserves the utility for training IDS models?
@@
    How well do these methods maintain the original data’s statistical properties (fidelity) and integrity?
@@
    Which techniques effectively balance class distributions in the generated data?
@@
    Which GAN models are best for producing cyber security network traffic data? These questions guide the comprehensive benchmarking of both statistical and AI-based synthetic data generation techniques.
I-E Contribution
The contributions of this work are summarized below:
  1. 1.
    Comprehensive Comparison of Synthetic Data Generation Methods: This research uniquely compares three prominent synthetic data generation methods—statistical, classical AI, and generative AI—on cybersecurity network traffic datasets (NSLKDD, CICIDS17). This provides the first direct comparison of these methods in the cybersecurity field.
  2. 2.
    Filling the Knowledge Gap in Cybersecurity Data: The study addresses a gap in the literature by identifying the most suitable synthetic data generation method specifically for network traffic datasets, offering clarity in an area where no previous comparative study exists.
  3. 3.
    Practical Guidance for Researchers: By evaluating the performance of these methods, this research offers a clear recommendation on the best approach for generating synthetic cybersecurity data, enabling researchers to make informed decisions without having to test multiple methods.
  4. 4.
    Introduction of Mutual Information for Feature Selection: Using Mutual Information for feature selection in synthetic data generation is an innovative contribution, providing a new perspective that may improve feature selection strategies in future cybersecurity studies.
The source code of this experiment is publicly available on our GitHub Repository  [^1] https://github.com/AdenRajput/Comparative_Analysis.git.
@@ I-F Paper Organization
The remaining sections of this paper are organized as follows:
  • Section II provides an overview of Synthetic Tabular Data Generation Methods, their usage in tabular data applications in cybersecurity, and evaluation metrics.
  • Section III details the prominent standard non-AI and AI-based synthetic tabular data generation methods and their mathematical workflow.
  • Sections IV and V describe the mutual information and evaluation metrics used in this experiment.
  • Section VI discusses the experimental setup, including the datasets and model configurations.
  • Section VII presents our experiments’ results, highlighting each method’s performance across two datasets.
  • Finally, in Sections VIII and IX, we mention the potential areas for future work and development, focusing on critical aspects that require further investigation and improvement, and conclude with key highlights of this study.

### v2

# Synthetic Network Traffic Data Generation: A Comparative Study

@@ Abstract
The generation of synthetic network traffic data is essential for network security testing, machine learning model training, and performance analysis. However, existing methods for synthetic data generation differ significantly in their ability to maintain statistical fidelity, utility for classification tasks, and class balance. This study presents a comparative analysis of twelve synthetic network traffic data generation methods, encompassing non-AI (statistical), classical AI, and generative AI techniques. Using NSL-KDD and CIC-IDS2017 datasets, we evaluate the fidelity, utility, class balance, and scalability of these methods under standardized performance metrics. Results demonstrate that GAN-based models, particularly CTGAN and CopulaGAN, achieve superior fidelity and utility, making them ideal for high-quality synthetic data generation. Statistical methods such as SMOTE and Cluster Centroid effectively maintain class balance but fail to capture complex traffic structures. Meanwhile, diffusion models exhibit computational inefficiencies, limiting their scalability. Our findings provide a structured benchmarking framework for selecting the most suitable synthetic data generation techniques for network traffic analysis and cybersecurity applications.
@@
Synthetic tabular Data, Network traffic, NSL-KDD, CIC-IDS, Mutual information, Generative Adversarial Networks (GANs), Diffusion models, generative AI
@@ I Introduction
The increasing reliance on data-driven decision-making in networking and security has intensified the need for high-quality network traffic data. This data plays a fundamental role in various applications, including network performance analysis, security research, and the development of machine learning (ML) models [1]. However, collecting and utilizing real-world network traffic data presents significant challenges. Issues such as privacy concerns, limited availability, and the high cost of manual labeling restrict access to large, diverse, and representative datasets [2].
To overcome these limitations, synthetic network traffic data generation has emerged as a promising alternative [3, 4, 5, 6, 7]. Synthetic datasets can simulate network traffic patterns while mitigating ethical, legal, and confidentiality concerns, enabling researchers and practitioners to develop, evaluate, and deploy data-driven solutions in a more scalable manner [8]. Despite this potential, existing methods for synthetic network traffic generation vary widely in their effectiveness, and no comprehensive comparison exists to guide the selection of the most suitable approach for synthetic network traffic data generation.
I-A The Need for Synthetic Network Traffic Data
Synthetic network traffic data generation is important for several reasons:
@@ I-A The Need for Synthetic Network Traffic Data
    Privacy and Data Sensitivity: Due to regulatory restrictions (e.g., GDPR) and commercial confidentiality, real network traffic data is often inaccessible [9]. Synthetic data provides a privacy-preserving alternative without exposing sensitive user information [10].
@@
    Data Imbalance and Scarcity: Real-world network traffic datasets are highly imbalanced, with benign traffic dominating over malicious activity, which can limit the effectiveness of ML-based models [11]. Synthetic data can address this by generating balanced datasets for better model training and evaluation.
@@
    Scenario Diversity: Real-world datasets may not cover a broad range of traffic conditions. Synthetic data generation techniques allow for the creation of diverse traffic scenarios, including rare or unseen network behaviors, which are crucial for testing robust ML models [12, 13].
@@
    Scalability and Adaptability: Unlike real data collection, which requires extensive monitoring infrastructure, synthetic data can be generated on demand to match specific analytical or experimental needs, reducing reliance on costly and time-consuming data collection processes [14].
I-B Challenges in Synthetic Network Traffic Generation
Despite its advantages, generating high-quality synthetic network traffic data remains challenging:
@@ I-B Challenges in Synthetic Network Traffic Generation
    Ensuring Realism: Many synthetic data generation methods struggle to replicate the statistical and temporal properties of real network traffic, limiting their practical applicability [15].
@@
    Variability and Generalization: Some generative models suffer from mode collapse, producing limited traffic variations, which reduces their effectiveness in representing diverse network conditions [16, 17, 18].
@@
    Computational Complexity: State-of-the-art generative methods, such as Generative Adversarial Networks (GANs) and diffusion models, require significant computational resources, which can hinder their scalability for real-world deployment [19].
@@
    Class Imbalance: Many real-world network datasets exhibit severe class imbalances, where attack or rare traffic classes are underrepresented. Existing synthetic data generation techniques often fail to address this imbalance effectively, leading to biased models that struggle with anomaly detection and classification [20].
@@ I-C Motivation
Generating synthetic network traffic data plays a critical role in network analysis, security testing, and developing machine learning models. While various methods exist for generating synthetic traffic, ranging from non-AI (statistical techniques) to AI (classical and generative) approaches, their effectiveness varies significantly [21, 22]. Prior studies have explored these methods individually, yet their comparative strengths and weaknesses remain unclear.
One of the key limitations in existing research is the lack of a systematic evaluation of synthetic data generation techniques across different methodological categories. Most studies focus on a single class of methods without directly comparing their ability to generate high-fidelity and utility-preserving data.
Another challenge lies in handling class imbalance in network traffic datasets. Many real-world network environments exhibit highly skewed distributions, where benign traffic significantly outnumbers attack or anomaly instances. While synthetic data generation has been proposed as a solution, current methods often fail to maintain class balance without compromising data fidelity [23].
To address these limitations, this study systematically evaluates and compares multiple synthetic network traffic generation methods across non-AI (statistical) and AI ( classical and generative) approaches. By assessing their fidelity, utility, and class-balancing capabilities, this research provides valuable insights into selecting the most suitable synthetic data generation method for different network traffic applications.
@@ I-D Objectives
This study aims to systematically evaluate synthetic network traffic data generation methods by addressing the following research questions:
@@
    Fidelity and Realism: How well do different synthetic data generation techniques preserve real network traffic’s statistical properties and temporal dynamics?
@@
    Utility for Machine Learning Tasks: How effectively can synthetic data generated by different methods improve the performance of ML-based network traffic classification models?
@@
    Class Balance and Data Diversity: Which methods most effectively address class imbalance in network traffic datasets while maintaining statistical similarity?
@@
    Scalability and Computational Efficiency: What are the computational trade-offs associated with each method, and how do they impact scalability for large-scale simulations?
I-E Contributions
This study presents a systematic benchmarking of synthetic network traffic generation methods across statistical (non-AI) and AI (classical and generative) AI. Our key contributions are:
1) Comprehensive Comparative Analysis: We systematically evaluate twelve prominent synthetic data generation methods using standardized metrics, including fidelity, utility for machine learning models, class balance, and computational efficiency. This study provides a structured comparison under a unified experimental framework, addressing the lack of direct performance comparisons in prior work.
2) Empirical Insights into Trade-offs Between Fidelity, Utility, and Computational Cost: Our analysis reveals that generative AI models achieve superior fidelity but require higher computational resources and often struggle with class balance. In contrast, statistical methods effectively handle class imbalance but fail to preserve complex traffic patterns. These findings provide practical recommendations for selecting the most suitable method based on specific application needs.
3) Evaluation of Class Imbalance Handling in Synthetic Network Traffic Generation: We analyze how different methods address the issue of class imbalance, a persistent challenge in real-world network traffic datasets. Our results highlight the effectiveness of various techniques in generating balanced yet realistic traffic distributions, contributing to the development of more reliable synthetic datasets for machine learning applications.
The source code for this study is publicly available at https://github.com/AdenRajput/Comparative_Analysis.git.
@@ I-F Paper Organization
The rest of this paper is structured as follows:
  • Section II reviews existing synthetic data generation methods and highlights gaps in prior research.
  • Section III describes the non-AI (statistical) and AI (classical and generative) based synthetic tabular data generation methods.
  • Section IV introduces the evaluation metrics used to assess fidelity, utility, and class balance.
  • Section V presents the experimental setup, including datasets, pre-processing steps, and model configurations.
  • Section VI provides a detailed comparative analysis of results across different methods.
  • Section VII outlines future research directions and potential improvements in synthetic network traffic generation.
Figure 1: Methods for generating synthetic tabular data
TABLE I: Comparison of Synthetic Network Traffic Data Generation Methods
Category Method Key Features Strengths Weaknesses
Non-AI (Statistical) Random Oversampling (ROS) [24] Randomly duplicates minority class samples Simple, low computational cost Can lead to overfitting
Synthetic Minority Oversampling Technique (SMOTE) [25] Generates synthetic samples via interpolation Generalized decision boundaries for classifiers and reduce overfitting Doesn’t consider majority class
Adaptive Synthetic Sampling (ADASYN) [26] Adaptive sample generation focused on hard-to-learn examples Addresses class imbalance dynamically Can introduce noisy samples
Cluster Centroids (CC) [27] Reduces majority class by replacing samples with centroids Avoids data duplication May oversimplify and loss key information about data structure
Gaussian Mixture Models (GMM) [28] Models data distribution probabilistically Captures complex distributions Sensitive to parameter tuning
AI-Based (Classical + Generative) Bayesian Networks [29] Probabilistic graphical models Captures dependencies between features Requires expert knowledge for structure learning
Tabular Variational Autoencoder (TVAE) [30] Encode-decoder architecture for learning latent representations Learns complex distributions Computationally expensive
Tabular Diffusion Model (TabDDPM) [31] Uses iterative denoising to generate data Preserves feature dependencies Computationally intensive
Conditional Tabular GAN (CTGAN) [32] Conditional generator and mode-specific normalization Handles mixed data types effectively Requires extensive hyperparameter tuning
CopulaGAN [33] Uses copula transformations in GANs Preserves feature dependencies Less effective in high-dimensional data
Naive Bayes and Logistic Regression GAN (GANBLR++) [34] Enhances GAN training with Bayesian learning Generates both categorical and numerical data using a Dirichlet Mixture Mode Requires accurate numerical indices
Cascaded Tabular GAN (CasTGAN) [35] Cascade training mechanism for tabular data Preserves feature dependencies Computationally intensive